OpenStudy (anonymous):

What is the simplified form of the quantity 4 x squared minus 25 over the quantity 2 x plus 5 ?

OpenStudy (anonymous):

\[\frac{4x^2-25}{2x+5}\] factor the numerator, then cancel

OpenStudy (triciaal):

do you recognize the numerator as the difference of 2 squares? 4x^2 = (2x)^2 and 25 = (5)^2 follow the directions already posted